Calculate the quantity of H+ions present in 200 cm3 of a solution of nitric acid with a concentration of 0,25mol.dm -3
1
Expert's answer
2020-06-04T10:16:06-0400

Solution:

Nitric acid (HNO3) is a strong acid.

This means that HNO3 dissociates completely to H+(aq) and NO3- (aq) when it dissolves in water.

HNO3(aq) = H+(aq) + NO3-(aq)

Therefore,

C(HNO3) = [H+] = 0.25 mol dm-3 = 0.25 mol/L


Moles of H+ = [H+] × Volume of solution = (0.25 mol/L) × (0.2 L) = 0.05 mol

Number of H+ ions = Avogadro's number × Moles of H+

Number of H+ ions = 6.022×1023 × 0.05 = 0.30×1023 = 3.0×1022 ions


Answer: 3.0×1022 ions of H+.